# Pitch: AI-assisted Tagging of Deepfake Audio Calls using Challenge-Response

### v3

@@ Abstract.
The rise of AI voice-cloning technology, particularly audio Real-time Deepfakes (RTDFs), has intensified social engineering attacks by enabling real-time voice impersonation that bypasses conventional enrollment-based authentication. To address this, we propose Pitch, a robust challenge-response method to detect and tag interactive deepfake audio calls. We developed a comprehensive taxonomy of audio challenges based on the human auditory system, linguistics, and environmental factors, yielding 20 prospective challenges. These were tested against leading voice-cloning systems using a novel dataset comprising 18,600 original and 1.6 million deepfake samples from 100 users. Pitch’s prospective challenges enhanced machine detection capabilities to 88.7% AUROC score on the full unbalanced dataset, enabling us to shortlist 10 functional challenges that balance security and usability.
For human evaluation and subsequent analyses, we filtered a challenging, balanced subset. On this subset, human evaluators independently scored 72.6% accuracy, while machines achieved 87.7%. Acknowledging that call environments require higher human control, we aided call receivers in making decisions with them using machines. Our solution uses an early warning system to tag suspicious incoming calls as "Deepfake-likely." Contrary to prior findings, we discovered that integrating human intuition with machine precision offers complementary advantages. Our solution gave users maximum control and boosted detection accuracy to 84.5%. Evidenced by this jump in accuracy, Pitch demonstrated the potential for AI-assisted pre-screening in call verification processes, offering an adaptable and usable approach to combat real-time voice-cloning attacks. Code to reproduce and access data at https://github.com/mittalgovind/PITCH-Deepfakes
@@ 1. Introduction
Since 2019, the mainstream adoption of online interactions has made telephony social engineering attacks increasingly plausible. The four predominant types of phone scams, in increasing order of their targeted nature, are robocalls, replayed voices, automated voice chatbots, and real-time voice conversion. This work focuses on the latter, exploring the benefits of challenge-response mechanisms in mitigating the risks associated with real-time voice conversions.
@@
Motivation. The scale of phone scams has reached alarming proportions. In 2022 alone, robocall-based phone scams in the United States soared to an astonishing 50.5 billion calls. By 2024, scammers had escalated their tactics, using robocalls to impersonate high-profile figures such as U.S. President Biden (nbc, 2024) and the Mayor of New York (APNews, [n. d.]) for voter suppression and outreach. This concerning escalation prompted repeated warnings from the FBI (Internet Crime Complaint Center (2021), IC3), the FTC (NPR, 2023), and an eventual ban enforced by the FCC (Federal Communications Commission, 2024).
@@
A more sophisticated and targeted form of impersonation fraud leverages Real-time Deepfakes (RTDFs). Scammers utilize speech-to-speech (voice conversion or VC) or text-to-speech (TTS) generation tools to make deepfake calls that sound convincingly similar to a target’s voice. These real-time, interactive calls are more believable and, consequently, more persuasive than traditional scams.
@@
While individual scam cases involving the impersonation of loved ones’ voices have been reported (Yorker, 2024), they can often be mitigated through simple precautions such as calling back, not being reactive, and asking personal questions. However, the challenge of differentiating imposters from genuine callers is far more acute and voluminous for contact centers of market entities such as banks, government agencies, and insurance companies. These organizations must continue serving millions of calls daily while protecting their clients’ sensitive accounts from impersonators.
@@
Companies rely on various security protocols to detect impersonators, including speaker verification systems, liveness detection, and knowledge-based authentication. However, a targeted RTDF-based (live deepfake) impersonation can potentially bypass all such conventional protections (Guide, 2023; NYTimes, 2023; Online, 2023). The situation is even more dire for contact centers of organizations that lack the infrastructure to support customers’ enrollment in such protective protocols, such as insurance companies and caller-ID services like Google Voice.
@@
While numerous audio deepfake detection methods exist (Yi et al., 2023), they are often brittle and primarily designed for static content. These traditional deepfake detectors base their judgment on a limited set of features and assume a non-interactive scenario. As these methods work offline, they are inefficient against deepfake calls where near real-time performance is desirable.
@@
Taking this integration a step further, in scenarios where humans were uncertain, machines were allowed to take charge and make the final call. This blend of Human-AI contributions, at a ratio of 56:44, enhanced the overall detection accuracy to 84.5%, representing a 16.4% improvement over the human-only baseline (with machine taking full control standing at 87.7% accuracy). This approach underscores that while humans maintain their primary role in decision-making, strategic machine involvement can dramatically enhance the accuracy of detecting fake calls.
@@
A challenge-response exploration to detect real-time audio deepfakes, focusing on making it interpretable (using human-in-the-loop) and scalable (using automated methods), and extensible to more sophisticated challenges, without assuming any prior information on the suspected caller.
A tagging framework, Pitch, leverages the complementary streng -ths of humans and AI to detect deepfake calls and adapts based on prediction confidence.
A taxonomy of speech tasks comprising eight categories and 22 subcategories that are challenging for state-of-the-art voice converters, evaluated using ML models and extensive user studies to assess the efficacy and usability of these challenges.
An open-source high-quality, manually-verified dataset with 18,600 original and 1.6 million deepfake samples covering 20 diverse challenges, including code and user-study instruments.
@@ 2.1. Real-time Deepfake Generation
Text-to-Speech. Text-to-Speech systems facilitate the conversion of written text into audible speech, enabling automated vocalization of textual content. The process encompasses several stages: textual analysis, phonemic conversion, and speech synthesis. Contemporary TTS methodologies employ various techniques, ranging from concatenative synthesis to advanced neural network-based approaches (suno-ai, 2023; tor, 2023). The primary objective of these systems is to generate naturalistic speech output that accurately conveys prosodic and intonational nuances inherent in human speech patterns.
@@ 2.3. Problem Description
Distinct Concepts. It is important to differentiate our work from related but distinct concepts:
@@ 3. Related Work
Challenge-Response Systems. Authentication schemes utilizing challenge-response mechanisms are exemplified by the ubiquitous Captcha, which differentiates bots masquerading as humans on the internet. Audio-Captcha (Fanelle et al., 2020) extends this approach for people with visual impairments. The Gotcha system (Mittal et al., 2024) introduced a challenge-response approach for real-time video-only deepfake detection, analyzing eight video challenges. This study concluded that challenges lead to consistent and measurable degradation in detecting real-time deepfakes in muted videos, with a focus on interpretability.
D-Captcha (Yasur et al., 2023) developed a challenge-response system for audio calls, evaluating nine challenges and measuring detection performance based on realism, compliance, and identity against deepfake generators, with StarGANv2-VC (Li et al., 2021) being the most potent deepfake generator tested. Their research provides a vital proof-of-concept for a more systematic and robust approach, indicating the promising potential of challenge-response systems in detecting deepfake audio calls. D-Captcha++ (Nguyen-Le et al., 2024) is a follow-up that proposed an imperceptible adversarial attack on D-Captcha (Yasur et al., 2023) and employs projected gradient descent-based adversarial training to defend against it.
@@
Novelty of Our Approach. To the best of our knowledge, Yasur et al. (Yasur et al., 2023) is the only closely related work that explores a challenge-response mechanism for detecting audio deepfakes. Our work expands upon this foundation in several key ways:
We explore a substantially broader spectrum of audio challenges and construct a comprehensive taxonomy based on the human auditory system, linguistics, and environmental factors.
We evaluate 20 distinct audio tasks, tested against a state-of-the-art voice converter. Notably, this modern converter rendered 6 out of 9 challenges from Yasur et al. ineffective, underscoring the rapid evolution of deepfake technology.
We introduce a novel human-AI collaborative perspective, integrating human oversight with machine efficiency to enhance detection accuracy and interpretability.
We contribute a high-quality speech dataset that significantly surpasses previous efforts in scale and diversity:
5x increase in the number of speakers
2x more challenge types
7.5x more original audio files
100x more deepfake samples
Inclusion of on-the-move smartphone environments
Comprehensive demographic diversity
These advancements extend the scope of challenge-response mechanisms in the context of audio deepfake detection into computer supported cooperative work and facilitates future research.
@@ 4.1. Audio Challenges and their Taxonomy
Table 1 delineates our taxonomy of audio challenges, designed to stimulate variations in speech factors such as linguistics, tone of voice, background, and behavioral cues. The following describes each category and sub-category.
resources/evaluation/taxonomy_only
@@
Lingual challenges prompt individuals to alter their vocal characteristics in distinctive ways, exploiting the degrees of freedom in the human vocal system.
@@ 4.1.1. Vocal Distortions
These challenges manipulate mouth peripherals and vocal chord resonance:
Vocal Peripherals: Impeding the movement of teeth, lips, or tongue (e.g., speaking with a pen between teeth) restricts articulation, potentially hampering speech recognition in RTDFs (Zhang et al., 2022).
Whisper: Whispering, characterized by minimal vocal cord vibrations, creates a toneless output that may challenge deepfakes trained on normal speech.
Vocal Cavity: Modifying the vocal tract cavity (e.g., speaking while holding one’s nose) impacts sound resonance, potentially confounding RTDFs.
Waveform Challenges. These involve modifying key speech characteristics in the sound waveform:
Frequency: Varying voice pitch, from unusually low or high pitches to singing melodies, tests the waveform diversity an RTDF can handle (Chowdhury et al., 2022).
Amplitude: Modulating speaking volume offers a natural test for RTDF detection.
Temporal: Abnormal speech rates may challenge RTDFs trained on average speaking speeds.
Language and Articulation Challenges. These involve articulating uncommon language patterns:
Hesitation: Pronouncing difficult sequences or foreign words tests RTDFs’ ability to handle "guessing" phonetic patterns.
Mimicry: Users mimicking known accents may confound RTDFs trained on specific datasets.
Phonetics: Reproducing sounds not native to the speaker’s language (e.g., rolling Rs) challenges RTDFs’ adaptability.
Deception: Unexpected commands to cease speaking test reaction times, can expose RTDFs’ lack of human-like responsiveness.
@@
Emotion: Speaking with discernible emotions tests RTDFs’ ability to replicate nuanced expressions.
Phonology: Altering inflection, particularly at sentence endings, challenges RTDFs’ prosodic capabilities.
@@
These challenges rely on human-generated noises, environmental sounds, or unique speech patterns.
@@
Vocal Noises: Includes coughs, whistles, and loud breathing.
Non-vocal Noises: Involves producing intentional sounds like clapping, snapping fingers, or microphone manipulation.
Environmental Noises: Incorporates background sounds like bird songs or traffic noise.
Reverb: Creates feedback loops or reverberations during the call.
@@
Playback. This category involves playing recordings that may disrupt deepfake creation:
Speech: Subjects speak while a recorded voice is playing, potentially creating artifacts as RTDFs struggle with multiple speakers.
Music: Speaking over music introduces non-speech audio patterns, challenging the voice conversion process.
Passive Distortions (out-of-scope): They involve manipulating the caller’s voice or playing adversarial audio, are considered out-of-scope as they assume trust in the caller’s equipment.

### v4

@@ Abstract.
The rise of AI voice-cloning technology, particularly audio Real-time Deepfakes (RTDFs), has intensified social engineering attacks by enabling real-time voice impersonation that bypasses conventional enrollment-based authentication. This technology represents an existential threat to phone-based authentication systems, while total identity fraud losses reached $43 billion. Between 2022-23, deepfake attempts occured once every five minutes, with 26% of Americans reporting they have encountered deepfake scams and 9% falling victim. Unlike traditional robocalls, these personalized AI-generated voice attacks target high-value accounts and circumvent existing defensive measures, creating an urgent cybersecurity challenge. To address this, we propose Pitch, a robust challenge-response method to detect and tag interactive deepfake audio calls. We developed a comprehensive taxonomy of audio challenges based on the human auditory system, linguistics, and environmental factors, yielding 20 prospective challenges. Testing against leading voice-cloning systems using a novel dataset (18,600 original and 1.6 million deepfake samples from 100 users), Pitch’s challenges enhanced machine detection capabilities to 88.7% AUROC score, enabling us to identify 10 highly-effective challenges.
For human evaluation, we filtered a challenging, balanced subset on which human evaluators independently achieved 72.6% accuracy, while machines scored 87.7%. Recognizing that call environments require human control, we developed a novel human-AI collaborative system that tags suspicious calls as "Deepfake-likely." Contrary to prior findings, we discovered that integrating human intuition with machine precision offers complementary advantages, giving users maximum control while boosting detection accuracy to 84.5%. This significant improvement situates Pitch’s potential as an AI-assisted pre-screener for verifying calls, offering an adaptable approach to combat real-time voice-cloning attacks while maintaining human decision authority. Access code and dataset at govindm.me/pitch
@@ 1. Introduction
Motivation. The scale of phone scams has reached alarming proportions. In 2022 alone, robocalls in the United States soared to an astonishing 50.5 billion instances, many of them leading to scams. By 2024, scammers had escalated their tactics, using robocalls to impersonate high-profile figures such as U.S. President Biden (NBC News, 2024) and the Mayor of New York (APNews, 2023) for voter suppression and outreach. Truecaller reports that Americans still suffer a $25B annual loss from phone scams, impacting 21% of the adult public (Truecaller, 2024).
@@
A more sophisticated and targeted form of impersonation fraud leverages Real-time Deepfakes (RTDFs). Scammers utilize speech-to-speech (voice conversion or VC) or text-to-speech (TTS) generation tools to make live deepfake calls that sound convincingly similar to a target’s voice. These real-time, interactive calls are more believable and, consequently, more persuasive than traditional scams. Recent data from Entrust shows that deepfake volumes have increased by 3,000% between 2022-23 (Entrust, 2025), with attempts now occurring at a rate of one every five minutes.
The human inability to detect such sophisticated voice clones presents a critical vulnerability: only 27% of people feel confident they could identify if a call from a friend or loved one was real or AI-generated (McAfee, [n. d.]). These deepfakes represent an unprecedented evolution in social engineering attacks, exploiting deeply-rooted human cognitive biases around voice recognition and trust that traditional security awareness training cannot adequately address. This vulnerability has been exploited in high-profile cases, including a finance worker being convinced to remit $25 million by an impersonation of their chief financial officer during a group video call (CNN, 2024), a mother being extorted for $50,000 ransom after hearing what she believed was her daughter’s distressed voice (CNN, 2023), and an energy company being defrauded of $243,000 by an impersonation of their boss’ voice (Wall Street Journal, 2019). McAfee found that 31% of respondents had experienced some kind of AI voice scam, believing they were hearing a friend or loved one when they were actually hearing a voice clone (McAfee, [n. d.]).
@@
While individual scam cases involving the impersonation of loved ones can often be mitigated through precautions such as calling back or asking personal questions, the challenge is far more acute for contact centers of financial institutions, government agencies, and insurance companies. These organizations must continue serving millions of calls daily while protecting their clients’ sensitive accounts from increasingly sophisticated impersonators.
@@
Companies rely on various security protocols to detect impersonators, including speaker verification systems, liveness detection, and knowledge-based authentication. However, a targeted RTDF-based impersonation can potentially bypass all such conventional protections (KYC AML Guide, 2023; NYTimes, 2023; CSO Online, 2023). This represents a fundamentally new attack vector in the cybersecurity landscape, where a compromised voice identity can cascade into broader system infiltration, credential theft, and data breaches that traditional security frameworks are not designed to counteract. The situation is even more dire for organizations that lack the infrastructure to support customers’ enrollment in such protective protocols.
@@
While numerous audio deepfake detection methods exist (Yi et al., 2023), they are often brittle and primarily designed for static content. These traditional deepfake detectors base their judgment on a limited set of features and assume a non-interactive scenario. As these methods work offline, they are inefficient against deepfake calls where near real-time performance is essential. With deepfakes now accounting for 40% of all biometric fraud attacks (Entrust, 2025), there is an urgent need for robust, real-time approaches to detect and mitigate audio deepfakes in interactive environments.
@@
Taking this integration a step further, in scenarios where machine predictions were highly confident, machines were allowed to take charge and make the final call. This blend of Human-AI contributions, at a ratio of 56:44, enhanced the overall detection accuracy to 84.5%, representing a 16.4% improvement over the human-only baseline (with machine taking full control standing at 87.7% accuracy). This approach underscores that while humans maintain their primary role in decision-making, strategic machine involvement can dramatically enhance the accuracy of detecting fake calls.
@@
A challenge-response system to detect real-time audio deepfakes that significantly advances prior work by providing stronger detection guarantees on a challenging dataset (57% relative AUC improvement versus 33%), focusing on making detection interpretable (using human-in-the-loop) and scalable (using automated methods), without assuming prior information on the caller.
A novel tagging framework, Pitch, that leverages the complementary strengths of human intuition and machine precision through confidence calibration, resulting in a 16.4% improvement in detection accuracy (84.5%) over the human-only baseline while maintaining human decision authority.
The first systematic and comprehensive taxonomy of speech tasks comprising eight categories and 22 subcategories based on auditory, linguistic, and environmental factors, with extensive evaluation confirming that 10 challenges consistently reveal deepfake vulnerabilities across multiple SoTA voice converters.
An open-source novel dataset that dramatically expands the scale of challenge-response evaluation with 18,600 original and 1.6 million deepfake samples (100× larger than prior work), including diverse recording environments, multiple devices, and demographically representative speakers.
@@ 2.1. Real-time Deepfake Generation
Text-to-Speech. Text-to-Speech systems facilitate the conversion of written text into audible speech, enabling automated vocalization of textual content. The process encompasses several stages: textual analysis, phonemic conversion, and speech synthesis. Contemporary TTS methodologies employ various techniques, ranging from concatenative synthesis to advanced neural network-based approaches (suno-ai, 2023; Betker, 2023). The primary objective of these systems is to generate naturalistic speech output that accurately conveys prosodic and intonational nuances inherent in human speech patterns.
@@ 2.2. Problem Description
Distinct Concepts. It is important to differentiate our work from these related concepts:
@@ 3. Related Work
Challenge-Response Systems. Authentication schemes utilizing challenge-response mechanisms are exemplified by the ubiquitous Captcha, which differentiates bots masquerading as humans on the internet. Audio-Captcha (Fanelle et al., 2020) extends this approach for people with visual impairments. The Gotcha system (Mittal et al., 2024) introduced a challenge-response approach for real-time video-only deepfake detection, analyzing eight video challenges. This study focused on interpretability and concluded that challenges lead to consistent and measurable degradation in detecting real-time deepfakes in muted videos.
| Feature | Pitch (Ours) | D-CAPTCHA (Yasur et al., 2023) | D-CAPTCHA++ (Nguyen-Le et al., 2024) | GOTCHA (Mittal et al., 2024) |
| Modality | Audio | Audio | Audio | Video |
| Key Innovation | Human-AI collaboration + confidence calibration | Challenge-based audio deepfake detection | Adversarial defense for D-CAPTCHA | Challenge-based video deepfake detection |
| Challenge Taxonomy | Extensive (8 categories, | |||
| 22 sub-categories) | Limited (no formal taxonomy) | Extends D-CAPTCHA | Facial Video-focused | |
| Number of Challenges | 20 evaluated, 10 effective | 9 evaluated, 3 effective | Built on D-CAPTCHA | 8 video challenges |
| Dataset Size | 18,600 original 1.6M deepfake 100 speakers | 2,500 original 16K deepfake 20 speakers | Based on D-CAPTCHA | Video only |
| Environment | Desktop & mobile, In-the-wild | Controlled laboratory | Controlled laboratory | Controlled laboratory |
| Deepfake Models Evaluated | FREE-VC StarGANv2-VC PPG-VC | StarGANv2-VC | StarGANv2-VC | N/A |
| Human-AI Integration | Collaborative & Calibration | None | None | Human-focused |
| Detection Performance | 88.7% AUC up from 56.0% (harder) | 95% Acc. up from 71% (easier) | Enhanced adversarial robustness | 88% Human, 80% Machine AUC |
D-Captcha (Yasur et al., 2023) developed a challenge-response system for audio calls, evaluating nine challenges and measuring detection performance based on realism, compliance, and identity against deepfake generators, with StarGANv2-VC (Li et al., 2021) being the most potent deepfake generator tested. Their research provides a vital proof-of-concept for a more systematic and robust approach, indicating the promising potential of challenge-response systems in detecting deepfake audio calls. D-Captcha++ (Nguyen-Le et al., 2024) is a follow-up that proposed an imperceptible adversarial attack on D-Captcha (Yasur et al., 2023) and employs projected gradient descent-based adversarial training to defend against it. Our work expands upon this foundation in several key ways as listed in Table 1, specifically by using our comprehensive taxonomy of audio challenges supporting human-AI collaborative integration on in-the-wild evaluations.
Comparison with Existing Authentication Protocols. Unlike traditional authentication methods (speaker verification, liveness detection, MFA, OTP), Pitch takes an active challenge-response approach to voice verification that addresses fundamental limitations in existing protocols.
Current speaker verification systems (Liu et al., 2024) require prior enrollment and struggle with natural voice changes over time, requiring regular re-enrollment and maintaining a secure infrastructure to maintain biometric information. Liveness detection mechanisms (Liu et al., 2018) are, in principle, ineffective against RTDFs, as the deepfake is already live and would be bypass liveness checks. Multi-factor authentication (MFA) introduces significant user friction and deployment complexities across communication channels, while one-time password (OTP) systems depend on secondary channels that are vulnerable to delays and real-time phishing during calls.
In contrast, Pitch requires no prior enrollment, trivially checks for liveness, operates entirely in-band during the call, while actively challenges potential impostors to perform tasks that exploit the fundamental limitations of speech synthesis technology. With 88.7% detection accuracy against SoTA deepfakes and a human-machine collaborative framework, Pitch provides a complementary security layer that offers superior interpretability through its challenge-response format. Rather than replacing existing authentication infrastructure, Pitch specifically addresses the emerging threat of real-time deepfakes while maintaining compatibility with other verification methods.
@@ 4.1. Audio Challenges and their Taxonomy
Table 2 delineates our taxonomy of audio challenges, designed to target fundamental limitations in current deepfake technology: (1) physiological modeling gaps in speech production, (2) limited training diversity in acoustic conditions, (3) inadequate handling of multi-source audio, (4) poor temporal adaptability in real-time processing, and (5) limited capability to maintain voice characteristics across diverse speech modes. The following describes each category and sub-category, with emphasis on how these challenges exploit specific weaknesses in deepfakes.
| Category | Sub-category | Examples (Selection Index) |
| No Challenge | Read Normally | Regular Speech (#0) |
| Vocal Peripherals | Static Mouth (#1), Cup Mouth (#2) | |
| Vocal Distortions | Whisper | Whispering (#3) |
| Vocal Cavity | Hold Nose (#4), humming | |
| Frequency | High (#5), Low pitch (#6), Sing (#7) | |
| Waveform | Amplitude | Speak Loudly (#8) & Softly (#9) |
| Temporal | Speak Quickly (#10) & Slowly (#11) | |
| Hesitation | Foreign Words (#12), Reverse Count | |
| Mimicry | Mimic another Accent (#13) | |
| Language / Articulation | Phonetics | Rolled R’s & Tongue clicks |
| Deception | Sudden Interruption while speaking | |
| Emotion | Sound happy / sad (#14) | |
| Tone of Voice | Phonology | Questions (inflection) (#15) |
| Vocal | Cough/whistle (#16) | |
| Non-vocal | Clap (#17), flick microphone | |
| Noise | Environmental | Birds, Cars |
| Reverb | Two Mics on same call | |
| Speech | Cross-talk (#18) | |
| Playback (desktop only) | Music | Instrumental (#19), Lyrical (#20) |
| Out-of-Scope | ||
| Unique Habits | Mannerisms, Person-of-Interest | |
| Behavioral | Biometric | Voice Print Detection |
| Perturbations | Adversarial Perturbations | |
| Passive Distortions | Software Editing | Modulating Pitch, Noise, or Bass |
| Lingual challenge, Non-lingual challenge, Replay challenge | ||
@@
Lingual challenges prompt individuals to alter their vocal characteristics in distinctive ways, exploiting the degrees of freedom in the human vocal system that are inadequately modeled by deepfakes.
@@ 4.1.1. Vocal Distortions
These challenges manipulate mouth peripherals and vocal chord resonance, targeting the deepfake weakness in physiological speech modeling:
Vocal Peripherals: Impeding the movement of teeth, lips, or tongue (e.g., speaking with a pen between teeth) restricts articulation. This exploits deepfakes’ inability to model the physical constraints of the human vocal tract, as current models lack explicit representation of articulatory dynamics (Zhang et al., 2022).
Whisper: Whispering, characterized by minimal vocal cord vibrations, creates a toneless output with significant spectral differences from normal speech. Deepfake models typically trained on normal speech samples lack sufficient exposure to this mode of vocalization, resulting in poor reconstruction quality.
Vocal Cavity: Modifying the vocal tract cavity (e.g., speaking while holding one’s nose) impacts sound resonance in ways that deepfake generators cannot accurately reproduce, as they do not incorporate acoustic principles of nasal and oral interaction.
Waveform Challenges. These involve modifying key speech characteristics in the sound waveform, exploiting the limited spectral and temporal diversity in deepfake training data:
Frequency: Varying voice pitch, from unusually low or high pitches to singing melodies, tests the waveform diversity an RTDF can handle. Most deepfake systems are trained on datasets with limited frequency range (e.g., VCTK corpus), resulting in degraded performance for outlier frequencies (Chowdhury et al., 2022).
Amplitude: Modulating speaking volume introduces variations in signal-to-noise ratio and harmonic structure that challenge the dynamic range capabilities of voice conversion systems, which are typically optimized for conversational amplitude levels.
Temporal: Abnormal speech rates stress deepfakes’ temporal modeling capabilities. Current systems struggle with maintaining linguistic coherence under extreme temporal constraints due to limitations in their sequence modeling architecture.
Language and Articulation Challenges. These involve articulating uncommon language patterns that exploit the limited linguistic diversity in training corpora:
Hesitation: Pronouncing difficult sequences or foreign words tests RTDFs’ ability to handle ‘guessing’ phonetic patterns, exposing their inability to model uncertainty in human conversations.
Mimicry: Users mimicking known accents may confound RTDFs as they typically assume consistent source speaker characteristics rather than intentional style manipulations.
Phonetics: Reproducing sounds not native to the speaker’s language (e.g., rolling Rs) challenges RTDFs’ adaptability to phonological structures absent in their training data.
Deception: Unexpected commands to cease speaking test reaction times, exposing RTDFs’ lack of human-like responsiveness and inability to model pragmatic aspects of communication.
@@
Emotion: Speaking with discernible emotions tests RTDFs’ ability to replicate nuanced expressions. Most voice converters focus primarily on content and speaker identity preservation, with emotional prosody often lost or flattened during conversion.
Phonology: Altering inflection, particularly at sentence endings, challenges RTDFs’ prosodic capabilities, as current models inadequately capture the full complexity of human intonation patterns, especially those that convey semantic nuance.
@@
These challenges rely on human-generated noises, environmental sounds, or unique speech patterns that exploit deepfakes’ limited ability to process complex acoustic environments.
@@
Vocal Noises: Includes coughs, whistles, and loud breathing. These non-speech vocalizations occupy spectral regions poorly modeled by deepfake systems, which are primarily optimized for linguistic content.
Non-vocal Noises: Involves producing intentional sounds like clapping, snapping fingers, or microphone manipulation. These create transient acoustic events that challenge voice conversion systems’ ability to maintain coherent voice characteristics during hybrid audio scenarios.
Environmental Noises: Incorporates background sounds like bird songs or traffic noise. Voice converters typically struggle with source separation when faced with complex acoustic environments, often degrading both the background and speech.
Reverb: Creates feedback loops or reverberations during the call, exploiting the limited ability of deepfakes to model acoustic environment properties and their effect on speech propagation.
@@
Playback. This category involves playing recordings that expose deepfakes’ weakness in handling multi-source audio:
Speech: Subjects speak while a recorded voice is playing, creating a multi-speaker scenario that forces voice conversion models to simultaneously process overlapping linguistic content, a capability not well-developed in current systems.
Music: Speaking over music introduces non-speech audio patterns with harmonic structures that interfere with the voice conversion process. Current deepfake systems lack robust source separation capabilities needed to cleanly process speech in such complex acoustic environments.
Passive Distortions (out-of-scope): These involve manipulating the caller’s voice or playing adversarial audio and are considered out-of-scope as they assume trust in the caller’s equipment.
By systematically exploiting these weaknesses, our challenges provide an effective framework for detecting audio deepfakes in real-world scenarios. To corroborate this, we begin by curating a data of audio challenges.